OpenStudy (anonymous):

I m sitll haveing problems with this problem. Not sure how to get this answer. The grade of a road is just the slope converted into a percent. So when you find the slope, multiply by 100 to get the grade. Again we use only non-negative slopes. A road rises 12 feet over a horizontal distance of 640 feet. What is its slope? What is its grade?

OpenStudy (anonymous):

\[tan \theta = 12/640\]

OpenStudy (anonymous):

So grade is equal to \[(12/640) \times 100\]

OpenStudy (anonymous):

1.875

OpenStudy (anonymous):

Is it clear

OpenStudy (anonymous):

Is it clear ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

Okay, bye then